OpenStudy (anonymous):

What is the approximate circumference of a circle with a radius of 40 inches? Use π ≈ 3.14. A. 125.6 inches B. 251.2 inches C. 502.4 inches D. 5024 inches i need help will fan and medal!

OpenStudy (texaschic101):

C = 2 * pi * r C = 2 * 3.14 * 40 C = 6.28 * 40 C = 252.2 inches any questions ?

OpenStudy (texaschic101):

oops..sorry....C = 251.2

OpenStudy (anonymous):

no questions thank you for explain it to me!

OpenStudy (texaschic101):

no problem :)